# CONLINE: Complex Code Generation and Refinement with Online Searching and Correctness Testing

### v1

@@ Abstract
Large Language Models (LLMs) have revolutionized code generation ability by converting natural language descriptions into executable code. However, generating complex code within real-world scenarios remains challenging due to intricate structures, subtle bugs, understanding of advanced data types, and lack of supplementary contents. To address these challenges, we introduce the CONLINE framework, which enhances code generation by incorporating planned online searches for information retrieval and automated correctness testing for iterative refinement. CONLINE also serializes the complex inputs and outputs to improve comprehension and generate test case to ensure the framework’s adaptability for real-world applications. CONLINE is validated through rigorous experiments on the DS-1000 and ClassEval datasets. It shows that CONLINE substantially improves the quality of complex code generation, highlighting its potential to enhance the practicality and reliability of LLMs in generating intricate code.
CONLINE: Complex Code Generation and Refinement
@@ 1 Introduction
Figure 1: An Example of the Human Programmer Code-writing Process Imitated by the CONLINE. After the problem is received, an online search is performed to simulate search results and create an initial version of code. Test cases are then generated, and the code is executed to produce output results. The code is refined based on the correctness of these results.
@@
Automated code generation from natural language descriptions has became more realistic, as large language models (LLMs) show their potential to generate complex and accurate code Manna and Waldinger (1971); Li et al. (2023); Luo et al. (2023); Rozière et al. (2024). There are a variety of methods that have been proposed to improve the quality of LLM code generation, such as retrieving offline documents Zhou et al. (2023); Jiang et al. (2023) and debugging generated code Zhang et al. (2023); Chen et al. (2023).
In real-world scenarios, complex code generation can involve intricate problem description, sophisticated code logic, and advanced data types Lai et al. (2022); Du et al. (2023); He et al. (2023). For example (see Figure 1), solving the required code generation involves multiple transformations of various matrices, as well as the invocation of functions that demands numerous and stringent parameters. Existing methods were not enough to address the following new challenges that emerge during complex code generation:
  1. 1.
    Complex code generation is hindered by intricate structures and logic. E.g., the complexity of the example in Figure 1 is as previously described.
  2. 2.
    Obscure bugs in complex code necessitate critical detection and refinement. Existing methods often refine code based on execution errors testing Zhang et al. (2023); Jiang et al. (2023), but for complex codes, sometimes the executable code has non-obvious bugs,e.g., initial code in Figure 1.
  3. 3.
    The complexity of inputs and outputs hampers LLM’s comprehension ability during code refinement. In complex code, inputs and outputs like Pandas DataFrames, PyTorch tensors, and matplotlib PNG images are intricate. Understanding such inputs and outputs poses challenges for LLMs due to large matrices, image modalities and so on.
  4. 4.
    Real-world scenarios frequently lack supplemental content, such as test cases and documentation pools. Existing works predominantly rely on pre-existing test cases in datasets Zhang et al. (2023); Jiang et al. (2023) or off-line document pools Zhou et al. (2023); Jiang et al. (2023), which are challenging to obtain directly in real-world scenarios.
@@
To address the above challenges, we propose a new code generation framework CONLINE that enhances the code generation and refinement quality of LLMs with planned online searching and automated correctness testing steps in this paper. The intuition of CONLINE is straightforward: During the process of coding, most human programmers are not bothered by the above challenges as shown in Figure 1. They could easily bypass them by searching on-line for answers, experiences, or guidelines. Also, they are able to design test cases and run code against them to verify the correctness of the code logic.
Correspondingly, CONLINE uses online search for code generation to partially address Challenges 1 and 4. For Challenge 1, it offers following advantages: (1) Retrieving information from library documents allows the retrieval of standardized function syntax, preventing errors from model hallucinations e.g., generating non-existing parameters. (2) Retrieving information from the latest blogs or Q&A platforms, such as StackOverflow, facilitates the emulation of commonly used code patterns, thereby reducing the complexity of generated code. For Challenge 4, online search encompasses the content of offline document pools and extends beyond with its superior capability to adapt to the diverse and evolving needs of real-world scenarios. Directly searching for solutions to such problems is difficult, because the problems are intricate and may involve multiple challenges. Therefore, we propose online search with generation query through planning.
To address the Challenge 2, this work proposes correctness testing in refinement. LLMs may generate complex code that contains subtle errors, resulting in code that is executable but does not produce the desired output. Therefore, during the refinement process, it is more critical to verify that the executed code produces the correct results rather than just checking for the existence of the errors. We propose incorporating both the execution output results and the errors as integral parts of the refinement to enhance the correctness of the code.
In refinement with correctness testing, we propose generation of test cases to address the Challenge 3. Test cases are crucial, as they serve as indispensable inputs for the code execution. This work utilizes LLMs to automatically generate test cases, making the framework applicable even in real-world scenarios where no known test cases exist. To address the Challenge 4, we propose serialization of input and output. It renders the input and output more intuitive and understandable for the model, particularly for those that are excessively long or non-textual modalities.
@@
We test the effectiveness of CONLINE on two complex code generation datasets using three different LLMs. For the complex code generation datasets, we selected DS-1000, which focuses on complex data science code generation, and the Class Eval dataset, which is oriented towards class-level code generation. By comparing with the existing state-of-the-art (SOTA) baseline, we achieved a 10.9% improvement on DS-1000 and average 9.47% on ClassEval, confirming the effectiveness of CONLINE. Through comprehensive ablation studies, we verified the effectiveness of each component of CONLINE. Moreover, given the varying performance of CONLINE on different base models, we discovered that CONLINE requires models to have capabilities such as Planning, which vary according to the complexity of the problem, in order to further enhance the model’s code generation abilities.
@@ 2 Related Work
The realm of automated code generation has been propelled by benchmark datasets such as HumanEval Chen et al. (2021), MBPP Austin et al. (2021), and APPS Hendrycks et al. (2021), which assess the proficiency of language models in generating executable code from descriptions. These datasets encompass a variety of programming problems, yet recent studies have sought to escalate the complexity of code generation tasks. Works like DS-1000 Lai et al. (2022), ClassEval Du et al. (2023) and Text2Analysis He et al. (2023) have introduced datasets targeting specialized domains, including data science, object-oriented class generation, and data analysis. These endeavors reflect an emerging trend towards enhancing models’ abilities to produce sophisticated and domain-specific code structures.
Recent research has concentrated on improving Large Language Models (LLMs) for code generation via retrieve, refine and so on strategies. Retrieve enhances model performance by pulling in external information, with techniques like DocPrompt Zhou et al. (2023) accessing fixed document repositories and SELVEVOLVE Jiang et al. (2023) using models as knowledge bases. However, these are limited to document-sourced content and provide limited benefits.
Refine iteratively enhances generated code for greater precision. Projects such as Self-Debug Chen et al. (2023) and SELFEVOLVE Jiang et al. (2023) focus on error refinement, while Self-EditZhang et al. (2023) uses code execution results as feedback. Despite advancements, these methods often miss subtle bugs and are less applicable to complex, real-world code generation without readily available test cases or ground truths.
Figure 2: The Pipline of CONLINE. Step 1: LLM is employed to strategize the Problem and formulate queries based on the outlined steps. These queries enable the retrieval of diverse information from the internet. A high-quality initial code can be obtained through effective planning and leveraging internet information. Step 2: LLM generates test cases for testing the initial code. The test results serve as crucial inputs for the subsequent cycle of code refinement. Through iterative refinement processes, the quality of the initial code can be significantly improved.
@@ 3 Methodology
In this work, we adopt a two-step approach for code generation, mirroring the way humans write code. The first step is retrieve, where relevant information is obtained through online saerch and utilized by LLMs to generate initial code. The second step is refine, where the initial code is refined based on the execution results, leading to the generation of the final version of the code.
3.1 Retrieve
The difficulty in achieving effective online retrieval lies in formulating optimal search queries. On the one hand, for complex code generation, the problems are intricate and may involve multiple challenges. Directly searching for solutions to such problems is inaccurate and difficult. On the other hand, it is challenging that matching queries directly through methods for offline document like similarity calculations, due to the nature of online libraries. Therefore, we propose using planning to solve the challenge.
The retrieval process is divided into three steps: 1. Search queries $Q=\{q_{1},\ldots,q_{N}\}$ are generated through planning. 2. Conducting online searches using these queries to obtain relevant background information $INFO=\{info_{1},\ldots,info_{M}\}$. 3. The initial code $W_{0}$ is generated by the LLMs $\theta$ with the information obtained $I$:
$\widehat{W}_{0}\sim p_{\theta}(.|D,I)$ (1)
@@ 3.1.2 Online Search
For the above generated queries, we conduct an online search. We use the online search API [^1] https://github.com/Nv7-GitHub/googlesearch for the search process.
$\{url_{1},\ldots,url_{N_{u}}\}=search(q_{j}),j\in\{1,\ldots,N_{q}\}$
where, $N_{q}$ is the number of queries for the problem, $N_{u}$ is the number of urls for one query. In this study, we use $N_{q}=1$ and $N_{u}=1$.
@@
3.2 Refine
Existing work Chen et al. (2023); Jiang et al. (2023) typically emphasizes correction of errors identified during the refinement process. However, we observe that refining code that produces error-free outputs is equally crucial during the refinement process. Therefore, we introduce correction testing in §3.2.1. Additionally, we propose methods for the generation of test cases and serialization of inputs and outputs during the refinement process.
@@ 3.2.1 Correctness Testing
$\left\{\begin{array}[]{ccl}e_{j,k},o_{j,k}&=&execute(W_{j},i_{k}),j\in\{1,\ldots,N_{f}\}\\ INFO_{e_{j,k}}&=&\{e_{j,k},extract(search(e_{j,k}))\}\\ \widehat{W}_{j+1}&\sim&p_{\theta}(.|D,W_{j},\{S_{i},S_{o_{j}},INFO_{e_{j}}\}_{k}),\\ &&\hfil k\in\{1,\ldots,N_{i}\}\end{array}\right.$
where, $N_{f}$ is the total number of refinement steps, $N_{i}$ is the number of inputs. $i_{K}$ is the k-th input for the problem from Equation (3), $S_{i}$ and $S_{o_{j}}$ is the serialization of the input and output from Equation (4).
3.2.2 Generation of Test Case
Test cases is crucial, as they serve as indispensable inputs for the code execution in refinement. While, existing works in refining code predominantly rely on pre-existing test cases in datasets Zhang et al. (2023); Jiang et al. (2023), which are challenging to obtain directly in real-world scenarios. Moreover, some existing work Chen et al. (2023) even uses the ground truth output of the test case to refine the code, which is even more challenging to obtain for complex code problems in real-world scenarios. Because their problems involve various logical operations, and even with input test cases, deriving answers directly without code-based computations is demanding.
In this study, we propose generation of test case with LLMs to adapt to real-world scenarios.
@@ 3.2.3 Serialization of Input and Output
    For NumPy arrays, Pandas DataFrames, PyTorch tensors, and TensorFlow tensors, the serialization includes a truncated string of data, data type, data shape, and statistical information.
@@
Table 1: Table of Main Results and Ablation Study for DS-1000. The base model for CONLINE is GPT-4. All metric numbers are represented as percentages. For each metric, the bold number indicates the highest performance.
@@ 4.1.1 Datasets
We conduct experience on two complex code generation datasets:
1. DS-1000 Lai et al. (2022): DS-1000 is a code generation benchmark with a thousand data science questions spanning seven Python libraries. The complexity of this dataset is manifested in two aspects. First, complexity arises from intricate logical reasoning required during code generation due to the complex nature of the problems. For example, on the DS-1000 dataset, the average length of problem descriptions is 140 words, whereas other commonly used code generation datasets such as HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) have lengths of 23 and 15.7 words, respectively. Secondly, the input-output involves various complex data structures related to data science, making the code logic intricate during transformations of the data.
2. ClassEval Du et al. (2023): ClassEval is the first class-level Python code generation benchmark designed to evaluate code generation models’ performance on a diverse set of object-oriented programming tasks. The dataset comprises a curated collection of 100 tasks. These tasks cover a wide range of concepts, including inheritance, polymorphism, encapsulation, etc. Each coding task is in the format of the class skeleton, outlining target method description inside the class. The complexity of this dataset resides in its abstraction and hierarchical class structure. Tested models must generate large-scale code units and establish connections between each target method within the entire class, rather than focusing solely on individual functions.
@@
4.1.2 Evaluation
  • DS-1000. we follow the original dataset using Pass@1 accuracy. This evaluation was conducted across total and four types of perturbations: Origin, Surface, Semantic, and Diff-Rewrite.
  • ClassEval. we follow the original dataset and evaluate model performance through Pass@$K$ metric. Specifically, we calculate both class-leval and method-level Pass@$K$ with $K=1,3,5$.
4.1.3 Base LLMs
This work primarily utilizes the GPT OpenAI (2023) series as the LLM base model to validate the effectiveness of the framework. GPT-4 is utilized in its gpt-4-32k-0613 version, while GPT-3.5 is utilized in the gpt-35-turbo-16k-0613 version. To further investigate the performance of CONLINE on both open-source and specialized code generation models, we have also employed WizardCoder Luo et al. (2023) as a base model with WizardCoder-Python-13B-V1.0 version.
4.1.4 Baselines
For the DS-1000, we selected three LLM-based frameworks as baselines.
1. DocPrompt Zhou et al. (2023): DocPrompting enhances the LLM by employing a fine-tuned retriever to fetch problem-relevant documentation from offline document pools. The model then conditions on these documents, along with the problem description, to generate code.
2. Self-Debugging Chen et al. (2023): This approach depends on a SQL application or Python interpreter to instruct language models in revising SQL commands or Python code containing errors. For the sake of a fair comparison, we utilize its "simple" variant.
3. SELFEVOLVE Jiang et al. (2023): Employs LLMs as both sources of knowledge and self-reflective programmers. During the self-reflective process, it refines the code by addressing bugs.
For the ClassEval, We select four LLM-based code generation models as baselines: Instruct-CodeGen [^2] https://huggingface.co/sahil2801/instruct-codegen-16B, SantaCoder Allal et al. (2023), Instruct-StarCoder [^3] https://huggingface.co/GeorgiaTechResearchInstitute/ starcoder-gpteacher-code-instruct , WizardCoder Luo et al. (2023)
It is important to note that the test cases involving the refinement process in the baselines mentioned above all use the test cases from the dataset designated for testing. However, within the context of the real-world scenario of CONLINE, test cases from the dataset should not be used within the framework. Without these test cases, they are entirely incapable of functioning, hence we have positioned our CONLINE at a disadvantage for comparison purposes. Despite this, we are still able to achieve SOTA performance.
4.2 Main Results
Regarding the DS-1000 dataset, the main results are shown in Table 1. CONLINE surpasses the current SOTA model, SELFEVOLVE, by 10.9%, establishing itself as the new SOTA. Especially under the Diff-Rewrite perturbation setting, CONLINE exceeds SELFEVOLVE by 19.95%, which demonstrates the effectiveness of CONLINE in generating complex code. CONLINE employs online search and correctness testing to allow the model to imitate existing code patterns, thereby reducing the difficulty of generating new code, and refines the details to further enhance the correctness of the code.
Table 2: Table of Main Results and Ablation Study for ClassEval. All metric numbers are represented as percentages. For each metric, the bold number indicates the highest performance.
@@ 4.2 Main Results
   GPT-4 36.2 39.3 43.5 58.6 64.9 67.3
CONLINE (GPT-3.5) 33.9 43.6 45.8 58.4 62.0 64.7
   w/o refinement of output 31.2 40.7 43.5 56.8 60.3 64.2
   w/o refinement of error 30.7 42.5 41.4 53.6 62.0 61.3
   w/o generation of test case 31.0 40.2 41.9 55.4 59.1 61.7
   w/o online retrieval 31.4 40.6 42.8 55.7 59.4 62.3
   GPT-3.5 27.0 32.1 35.4 52.3 54.6 59.4
For the ClassEval dataset, the results are shown in Table 2. Our experiments demonstrate that CONLINE has an overall higher porformance on both class-level and method-level Pass@$K$ evaluation. Specifically, COLINE(GPT-4) outperforms the WizardCoder (best baseline model) significantly by an average of 31.3% on Class-level and 28.3% in Method level; COLINE(GPT-3.5) outporforms the WizardCoder by an average of 22.7& in Class-level and 17.2% in Method-level.

### v2

# CoCoST: Automatic Complex Code Generation with Online Searching and Correctness Testing

@@ Abstract
Large Language Models have revolutionized code generation ability by converting natural language descriptions into executable code. However, generating complex code within real-world scenarios remains challenging due to intricate structures, subtle bugs, understanding of advanced data types, and lack of supplementary contents. To address these challenges, we introduce the CoCoST framework, which enhances complex code generation by online searching for more information with planned queries and correctness testing for code refinement. Moreover, CoCoST serializes the complex inputs and outputs to improve comprehension and generates test cases to ensure the adaptability for real-world applications. CoCoST is validated through rigorous experiments on the DS-1000 and ClassEval datasets. Experimental results show that CoCoST substantially improves the quality of complex code generation, highlighting its potential to enhance the practicality of LLMs in generating complex code.
CoCoST: Automatic Complex Code Generation
@@ 1 Introduction
Figure 1: An Example of the Human Developer Code-writing Process Imitated by the CoCoST. After the problem is received, an online search is performed to simulate search results and create an initial version of code. Test cases are then generated, and the code is executed to produce output results. The code is refined based on the correctness of these results.
@@
Automatic code generation from natural language descriptions is becoming more realistic, as large language models (LLMs) show their potential to generate accurate code Li et al. (2023); Luo et al. (2023); Rozière et al. (2024). Various methods have been proposed to improve the quality of LLM code generation, such as retrieving offline documents Zhou et al. (2023); Jiang et al. (2023) and debugging generated code Zhang et al. (2023); Chen et al. (2023). However, complex code generation is a more difficult task, which involves intricate problem description, sophisticated code logic, and advanced data types Lai et al. (2022); Du et al. (2023); He et al. (2023). The existing methods struggle to address the arising challenges:
Challenge 1: Offline documents cannot meet the diverse demands of code generation. In real-world scenarios, these demands often exceed the capabilities of limited offline documents. For example, problem descriptions may involve functions that are not covered by pre-collected documents. Additionally, complex code generation for diverse needs often entails highly complex logic and a series of transformation functions like the programming problem in Figure 1, where simple API examples in documents fail to provide adequate guidance.
Challenge 2: In real-world situations, there is often a shortage of test cases (e.g., test cases in Figure 1) for automatic code generation. Most existing work depends heavily on pre-existing test cases in datasets Zhang et al. (2023); Jiang et al. (2023) , which are difficult to acquire directly in practical scenarios.
Challenge 3: Hidden bugs in complex code require meticulous identification and refinement. Current techniques frequently enhance code by analyzing execution errors Zhang et al. (2023); Jiang et al. (2023). But in the case of complex code, the executable code sometimes contains hidden bugs like the highlighted part of the initial code in Figure 1.
@@
To address these challenges, we introduce a new code generation framework named CoCoST [^1] The code will be publicly available upon acceptance. (Automatic Complex Code Generation with Online Searching and Correctness Testing) that improves the generation and refinement of complex code by LLMs through the planned online searching and automatic correctness testing steps. The intuition of CoCoST is straightforward: During the coding process, most human developers are not bothered by the above challenges, as illustrated in Figure 1. Developers can easily overcome these obstacles by searching online through engines (e.g., Google and Bing) for solutions, experiences, and guidelines. In addition, they can create test cases and execute code to ensure the correctness of the code logic.
To address Challenges 1, CoCoST proposes an online search methodology. This process involves querying web search engines and then extracting pertinent information to construct prompts for LLMs. The approach presents several benefits: (1) Retrieving information from the up-to-date blogs or Q&A platforms, such as StackOverflow, facilitates the emulation of commonly used code patterns, thereby reducing the complexity of generated code. (2) Online search extends beyond the scope of static offline documentation, covering a wider array of problems without being confined to a predetermined set. Meanwhile, it reduces the effort developers need to expend in assembling documentation, thereby increasing the framework’s level of automation. However, using problem descriptions as search queries can be challenging due to the complexity of the problems, which often encompass multiple issues. Therefore, we propose an online search with generation query through planning.
To address the Challenge 2, we introduce generation of test cases during refinement. Several studies Chen et al. (2022); Shinn et al. (2023) have attempted to generate tests. However, these methods often fall short when applied to the generation of complex code due to its intricate logic and outputs, which complicate the direct production of accurate tests (both the inputs and expected outputs for the solution code). CoCoST utilizes LLMs to automatically generate test cases (the inputs for the code). This strategy cleverly focuses on generating test cases without attempting to produce complete tests. It significantly simplifies the process of test case generation and facilitates its precise creation for complex code.
To address the Challenge 3, this work prioritizes correctness testing in refinement. During the refinement process, it is more critical to verify that the executed code produces the correct results rather than just checking for the existence of the errors. CoCoST incorporate both the execution output results and the errors into within the refinement prompts for LLMs to enhance the correctness. Moreover, during refinement, sophisticated data types and structures (within complex code itself, its inputs, and its execution results) are challenging for LLMs to understand, e.g., large Pandas DataFrames, Matplotlib charts. Thus, CoCoST propose serialization of input and output to convert them into understandable sequences before being processed by LLMs. Particularly those are excessively long or non-textual modalities.
@@
We evaluate the effectiveness of CoCoST on two complex code generation datasets (DS-1000 and ClassEval). By comparing with the existing state-of-the-art (SOTA) baseline, we achieve a 7.8% improvement on DS-1000 and average 9.47% on ClassEval.Moreover, we analyze and discovery that CoCoST requires models to have different capabilities such as planning, which vary according to the complexity of the problem.
In summary, our main contributions are:
  • We propose the novel CoCoST framework to generate complex code, and can be automatic in real-world scenarios.
  • We apply online search in code generation for the first time to our knowledge, in order to generate the complex code.
  • We prioritize correctness testing in refinement with test case generation and serialization of input and output, in order to refine hidden bugs in complex code.
  • We conduct experiments on DS-1000 and ClassEval datasets anddemonstrate the effectiveness and universality of CoCoST.
@@ 2 Related Work
Code generation datasets. The realm of automated code generation has been propelled by benchmark datasets such as HumanEval Chen et al. (2021), MBPP Austin et al. (2021), and APPS Hendrycks et al. (2021), which assess the proficiency of language models in generating executable code from descriptions. These datasets encompass a variety of programming problems, yet recent studies have sought to escalate the complexity of code generation tasks. Works like DS-1000 Lai et al. (2022), ClassEval Du et al. (2023) and Text2Analysis He et al. (2023) have introduced datasets targeting specialized domains, including data science, object-oriented class generation, and data analysis. These endeavors reflect an emerging trend towards enhancing models’ abilities to produce sophisticated and domain-specific code structures. In this paper, we select datasets with complex code generation to evaluate CoCoST.
Retrieval-augmented code generation. With the emergence of Large Language Models (LLMs), a variety of retrieval-augmented techniques have been developed to compensate for issues such as the inherent knowledge limitations. DocPromptZhou et al. (2023) and SELVEVOLVEJiang et al. (2023) leverage document libraries or models as knowledge bases to improve code generation. However, their reliance on fixed document libraries limits the scope of information they can provide and confines the generated code to the context of these libraries. Furthermore, the prerequisite of preestablished document libraries prevents these approaches from being fully autonomous in real-world frameworks. Solutions such as WebGPT Nakano et al. (2022), LaMDA Thoppilan et al. (2022), and FreshLLMs Vu et al. (2023) enhance the performance of natural language tasks by using online search or open web knowledge. However, because complex code generation often involves multiple steps and complexities, these methods struggle with direct application to complex code generation.
Code refinement. Refine iteratively enhances generated code for greater precision. Self-DebugChen et al. (2023), SELFEVOLVEJiang et al. (2023), and Self-EditZhang et al. (2023) improve code generation by refining code through the resolution of errors identified during execution. These methods effectively address errors, while when it comes to complex code generation, subtle bugs also play a significant role in the overall error landscape. Moreover, relying on pre-existing tests from datasets in refinement limits their autonomy in real-world applications, where such tests may not be readily available. CodeT Chen et al. (2022), Reflexion Shinn et al. (2023), and CODECHAIN Le et al. (2023) seek to strengthen code generation by creating tests. But the tests they generate include not only the inputs for the solution code but also the expected outputs. This poses a substantial challenge for complex code generation, where the logic can be intricate and certain problems may not lend to straightforward ground truth generation.
Figure 2: The Pipline of CoCoST. Step 1: LLM is employed to strategize the Problem and formulate queries based on the outlined steps. These queries enable the retrieval of diverse information from the internet. A high-quality initial code can be obtained through effective planning and leveraging internet information. Step 2: LLM generates test cases for testing the initial code. The test results serve as crucial inputs for the subsequent cycle of code refinement. Through iterative refinement processes, the quality of the initial code can be significantly improved.
@@ 3 Methodology
In this work, we adopt a two-step approach for code generation, mirroring the way humans write code. The first step is retrieval, where relevant information is obtained through online search and utilized by LLMs to generate initial code. The second step is refinement, where the initial code is refined based on the execution results, leading to the generation of the final version of the code.
3.1 Retrieval
The difficulty in achieving effective online retrieval lies in formulating optimal search queries. On the one hand, for complex code generation, the problems are intricate and may involve multiple challenges. Directly searching for solutions to such problems is inaccurate and difficult. On the other hand, it is challenging that match queries directly through methods for offline documents like similarity calculations, due to the nature of online libraries. So we propose generating queries through planning to solve the challenge.
The retrieval process is divided into three steps: 1. Search queries $Q=\{q_{1},\ldots,q_{N}\}$ are generated through planning. 2. Conducting online searches using these queries to obtain relevant background information $INFO=\{info_{1},\ldots,info_{M}\}$. 3. The initial code $W_{0}$ is generated by the LLMs $\theta$ with the information obtained $INFO$:
$\widehat{W}_{0}\sim p_{\theta}(.|D,INFO)$ (1)
@@ 3.1.2 Online Search
For the above generated queries, we conduct an online search. In this study, we use the online search API [^2] https://github.com/Nv7-GitHub/googlesearch for the search process as Equation (3). CoCoST can also be applied to private or domain-specific knowledge repositories as long as they are accessible via query, with details in §A.
$\{url_{1},\ldots,url_{N_{u}}\}=search(q_{j}),j\in\{1,\ldots,N_{q}\}$ (3)
where, $N_{q}$ is the number of queries for the problem, $N_{u}$ is the number of urls for one query. In this study, we use $N_{q}=1$ , $N_{u}=1$.
@@
3.2 Refinement
Existing work Chen et al. (2023); Jiang et al. (2023) typically emphasizes the correctness of errors identified during the refinement process. However, we observe that refining code that produces error-free outputs is equally crucial during the refinement process. Therefore, we introduce correctness testing in §3.2.1. Additionally, we propose methods for the generation of test cases and serialization of inputs and outputs during the refinement process.
@@ 3.2.1 Correctness Testing
[…]
where, $N_{f}$ is the total number of refinement steps, $N_{i}$ is the number of inputs. $i_{K}$ is the k-th input for the problem from Equation (4), $S_{i}$ and $S_{o_{j}}$ is the serialization of input and output from Equation (5).
3.2.2 Generation of Test Cases
Test cases are crucial, as they serve as indispensable inputs for the code execution in refinement. While, existing works in refining code predominantly rely on pre-existing test cases in datasets Zhang et al. (2023); Jiang et al. (2023), which are challenging to obtain directly in real-world scenarios. Moreover, some existing work Chen et al. (2023) even uses the ground truth output of the test case to refine the code, which is even more challenging to obtain for complex code problems in real-world scenarios. Because their problems involve various logical operations, deriving answers directly without code-based computations is demanding.
CoCoST introduces generation of test cases with LLMs to adapt to real-world scenarios.
@@ 3.2.3 Serialization of Input and Output
    For NumPy arrays, Pandas DataFrames, PyTorch tensors, and TensorFlow tensors, the serialization includes data truncated string, data type, data shape, and statistical information.
@@
Table 1: Main Results and Ablation Study for DS-1000. The base model for CoCoST is GPT-4. All metrics are represented as percentages. For each metric, the bold number indicates the highest performance.
@@ 4.1.1 Datasets
We conduct experiences on two complex code generation datasets:
DS-1000 Lai et al. (2022): DS-1000 is a code generation benchmark with a thousand data science questions spanning seven Python libraries. The complexity of this dataset is manifested in two aspects. First, complexity arises from intricate logical reasoning required during code generation due to the complex nature of the problems. For example, on the DS-1000 dataset, the average length of problem descriptions is 140 words, whereas other commonly used code generation datasets such as HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) have lengths of 23 and 15.7 words, respectively. Secondly, the input-output involves various complex data structures related to data science, making the code logic intricate during transformations of the data. Further details of DS-1000 implementation are shown in §B.1.
ClassEval Du et al. (2023): ClassEval is the first class-level Python code generation benchmark designed to evaluate code generation models’ performance on a diverse set of object-oriented programming tasks. The dataset comprises a curated collection of 100 tasks. These tasks cover a wide range of concepts, including inheritance, polymorphism, encapsulation, etc. Each coding task is in the format of the class skeleton, outlining the target method description inside the class. The complexity of this dataset resides in its abstraction and hierarchical class structure. Tested models must generate large-scale code units and establish connections between each target method within the entire class, rather than focusing solely on individual functions.
@@
Table 2: Main Results and Ablation Study for ClassEval. All metric numbers are represented as percentages. For each metric, the bold number indicates the highest performance.
@@
     w/o all (GPT-4 only) 36.2 39.3 43.5 58.6 64.9 67.3
4.1.2 Evaluation
DS-1000. We follow the original dataset using Pass@1 accuracy. This evaluation is conducted across total and perturbations: Origin, Surface, Semantic, and Diff-Rewrite.
ClassEval. We follow the original dataset using Pass@$K$ metric. We calculate both class-leval and method-level Pass@$K$ with $K=1,3,5$.
4.1.3 Base LLMs
This work primarily utilizes the GPT OpenAI (2023) series as the LLM base model to validate the effectiveness of the framework. GPT-4 is utilized in gpt-4-32k-0613 version, while GPT-3.5 is utilized in the gpt-35-turbo-16k-0613 version. To further investigate the performance of CoCoST on both open-source and specialized code generation models, we have also employed WizardCoder Luo et al. (2023) as a base model with WizardCoder-Python-13B-V1.0 version.
4.1.4 Baselines
For the DS-1000, we selected four LLM-based frameworks as baselines: DocPrompt Zhou et al. (2023), Self-Debugging Chen et al. (2023), SELFEVOLVE Jiang et al. (2023) and Reflexion Shinn et al. (2023). DocPrompting enhances the LLM by employing a fine-tuned retriever to fetch problem-relevant documentation from offline document pools. Self-Debugging depends on a Python interpreter to instruct language models in revising Python code containing errors. SELFEVOLVE employs LLMs as both sources of knowledge and self-reflective programmers. Reflexion utilizes reflective feedback with generated tests and episodic memory to process task feedback. Details are shown in §B.3.
For the ClassEval, we select five LLM-based code generation models and frameworks as baselines: Instruct-CodeGen [^3] https://huggingface.co/sahil2801/instruct-codegen-16B, SantaCoder Allal et al. (2023), Instruct-StarCoder [^4] https://huggingface.co/GeorgiaTechResearchInstitute/ starcoder-gpteacher-code-instruct , WizardCoder Luo et al. (2023) and Reflexion Shinn et al. (2023).
4.2 Main Results
Regarding the DS-1000 dataset, the main results are shown in Table 1. CoCoST surpasses the current SOTA framework, SELFEVOLVE, by 10.9%, establishing itself as the new SOTA. Especially under the Diff-Rewrite perturbation setting, CoCoST exceeds SELFEVOLVE by 19.95%, which demonstrates the effectiveness of CoCoST in generating complex code. CoCoST employs online search and correctness testing to allow the model to imitate existing code patterns, thereby reducing the difficulty of generating new code, and refining the details to further enhance the correctness of the code.
For the ClassEval dataset, the results are shown in Table 2. Our experiments demonstrate that CoCoST has an overall higher performance on both class-level and method-level Pass@$K$ evaluation. Specifically, CoCoST outperforms the Reflexion (best baseline model) significantly by an average of 19.5% and 20.4% on the Class and Method level.